Expand (ax-/xQj9 fine the oth terry Iconstant & summation of a ce terms in
Answer #1

де. have Bionomial expension, (x + y) 2 . m) xk ynak k=0 (9-) 9 (2x- :) ! (9-3) (23-%) = 3 62769* (**0 (9) (2x3° Estas) +Hence (14 term ан ереоріч , 5 9-r S (2) ( - ) (2) *. *.** ( ) * * * , — ) с 9! S х 2 х - 3 5! В-5)! 9. 3. 1.4 (П -3 х, 5 x 2